THAICONV(1) THAICONV(1) NAME thaiconv−converttextstringsinTIS620formatintoaspecialformattobeusedwithemacs20 incooperationwithcjk-enc.el. Itisnotintendedforstand-aloneuse. ThisfilterispartoftheCJKmacropackageforLATEX2ε. SEEALSO theCJKdocumentationfiles. AUTHOR WernerLemberg<wl@gnu.org> CJKVersion4.3.0 20-Jun-1999 1